OpenStudy (anonymous):

Jeannie is 8years older than her younger sister April. When added together, their ages equal 196.how old is April?

OpenStudy (goformit100):

let April's age be x. then Jeannie will be x+8.

OpenStudy (goformit100):

when added together means :- x + x+8 = 196 or, 2x+8 = 196 or, 2x + 188 or, x= 188/2 or, x= ?